Repeater usage guide v. 0.93

Greg Troxel, N1DAM

August 18, 1993

Contents

1 General 1

2 Repeater behavior 1 2.1 Frequencies, modulation . . . . . . . . 2 2.2 Location, antenna, power levels …. 2 2.3 Identifier behavior . . . . . . . . . . . 2 2.4 Tail . . . . . . . . . . . . . . . . . . . 2 2.5 Time-out timer . . . . . . . . . . . . . 2 2.6 Touch tones . . . . . . . . . . . . . . . 2

3 Timing, audio level, etc. 2

4 Identification 3

5 Listening, calling and conversation 3 5.1 Phonetics . . . . . . . . . . . . . . . . 3 5.2 Listening . . . . . . . . . . . . . . . . 4 5.3 Calling . . . . . . . . . . . . . . . . . . 4 5.4 Answering . . . . . . . . . . . . . . . . 5 5.5 Fine points about calling . . . . . . . . 5 5.6 Conversation . . . . . . . . . . . . . . 5 5.7 Making a call while a QSO is in progress 5 5.8 Joining an existing QSO . . . . . . . . 6

6 Signal reports 6 6.1 Audio . . . . . . . . . . . . . . . . . . 6 6.2 RF . . . . . . . . . . . . . . . . . . . . 7 6.3 Which RF path is the problem? …. 7 6.4 Signal strength meters: what they

don’t mean . . . . . . . . . . . . . . . 7 6.5 Giving reports . . . . . . . . . . . . . 7 6.6 Understanding your signal strength . . 8

7 Simplex 8

8 Touch tones 9 9 Emergency traffic 9

9.1 Asking for help . . . . . . . . . . . . . 9 9.2 Giving help . . . . . . . . . . . . . . . 9 9.3 Other aspects of emergency traffic . . 9

1 General

This is my attempt to explain to new operators what is common practice among experienced operators on the MIT repeater.

Many new people have asked me questions about how the repeater behaves or about operating practices, and I have written this document in an attempt to answer all of them in one place. It also contains some information of interest to experienced hams new to the MIT repeater, as it explains about the time-out timer, ID behavior and other repeater parameters.

The document is not finished; comments in italics that don’t make sense are notes to the author to look up and insert more info. Constructive criticism is invited; many changes have already been made as a result of comments of earlier readers. Please email

comments to gdt@n1dam.lcs.mit.edu.

This guide assumes you have studied for and passed the exams for a Technician class amateur radio license. Occasionally some technical statements are not explained; the guide does not attempt to be a tutorial about anything besides local conventions of repeater operating.

As a holder of an FCC amateur radio license, you should have read and have a copy of 47 CFR 97, FCC rules governing amateur radio. This is often referred to as “Part 97”, and can be purchased from several sources. A copy is

available via ftp on gandalf.umcs.maine.edu as
/pub/ham-radio/part97.txt, but it is missing Sub-

part E, prohibited practices. The ARRL sells a book with the rules and extensive interpretation for $9, and W5YI sells a booklet with mostly just the rules for $3. These should be available at most ham stores.^{1}

1ARRL: 225 Main St., Newington, CT 06111; W5YI: Fred-

The repeater operates under the amateur license of one of the club members, called the trustee or licensee. A repeater operating under a club callsign has a trustee, who holds the license on behalf of the club. A repeater operated by a club without a club callsign normally designates a member to serve as the “licensee”, and the repeater identifies with that member’s callsign.

When a person makes a transmission that is repeated, both the person transmitting and the trustee are responsible for the transmissions. The trustee has designated several people “control operators”. According to 97.103.a, these people are responsible for ensuring that transmissions from the repeater are in compliance with the rules. These people tend to listen to the repeater, and can turn it on and off as necessary.

The current trustee is N1CPK (Bill). Control operators are N1DAM (Greg), WA2EYC (Ron), and Steve (W1GSL).

The MIT repeater is different in character than many other repeaters, particularly the more popular 2m machines. Many people listen most of the time, but don’t always transmit. I, for example, usually have a receiver on when awake; other people also listen a lot but do not transmit as frequently. Thus, if you “call listening” (see below) and no one answers, that does not mean that no one heard you or that no one is listening. However, if you need help, say so — people who are listening but busy will almost always respond to calls for help even if they are too busy to talk.

2 Repeater behavior

This section attempts to explain how the repeater functions. It is somewhat detailed, and is intended not to be simplified further than is possible; thus it should not contain any misstatements. Some of this may not make sense; I do not attempt to explain the technical terms used, as such explanations are available in the literature. It does not attempt to explain how you should use the repeater; this is covered in later sections.

2.1 Frequencies, modulation

The repeater transmits on a nominal frequency 449.725 MHz using FM modulation with approximately 5 kHz deviation, resulting in a bandwidth of approximately 20 kHz. The transmit frequency should be within 1 kHz of nominal. The repeater receiver listens on 449.725 MHz. The receiver is FM with the same modulation parameters as the transmitter. Again, the receive frequency may differ slightly from what it should be, but should be within 1 khz.

The receiver operates in “tone squelch”, often called CTCSS or PL. In order for the receiver squelch to open, a tone of 114.8 Hz (PL code 2A) must be present on the incoming signal.

The repeater always transmits a CTCSS (or PL) tone. The frequency is 114.8 Hz.

2.2 Location, antenna, power levels

If you don’t understand part of this, don’t worry. The repeater is located at the top of MIT building 54, the Green building. The building is located at approximately 42 21 37 N, 71 05 24 W, or Maidenhead grid square FN42KI96. The antenna height is approximately 120 m. The antenna is vertically polarized, omnidirectional, and has gain of approximately 10 dBi. The repeater transmitter delivers approximately 50 W to the antenna. The repeater receiver has a noise figure of approximately 3 dB, allowing for the preamp, feedline and duplexor.

2.3 Identifier behavior

While in use, the repeater periodically transmits “W1XM/R” at 20 wpm in Morse code. Sometimes the sequence “W1XM/R MIT” is sent instead. Either of these is referred to as the “ID”. The repeater identifier has 2 basic states, idle and active. If the repeater has not transmitted for greater than 5 minutes, it is in the idle state. If the repeater is idle, and a transmission is received, the repeater will send “W1XM/R MIT” (the initial ID) at the end of the transmission. The repeater is then in the active state. If between 4 minutes 30 seconds and 5 minutes after the sending of an ID, a received transmission ends, the repeater sends “W1XM/R” (the normal ID). If 5 minutes pass since the sending of an ID, and there has been a transmission at any time during those 5 minutes, the repeater sends the initial ID. If there have not been any transmissions in the 5 minutes since the last ID, the repeater changes to the idle state.

erick O. Maia, W5YI, PO Box 565101, Dallas, Texas 75356; Ham Radio Outlet: 224 No. Broadway, Salem NH

2.4 Tail

The repeater has a 2 second “tail”. After a transmission that is received and retransmitted ends, the transmitter remains on for 2 seconds without any modulation.

Thus, after hearing a person’s transmission, there will be a burst of high-frequency noise at the end of the transmission, followed by a short period of silence, followed by another burst of high-frequency noise. The short burst of high-frequency noise is called a “squelch tail” because this is the open received channel noise being retransmitted before the squelch has a chance to close. The short period of unmodulated carrier is also called a “tail”, but is not a “squelch tail”.

By listening carefully, you can determine whether the repeater is still transmitting or not by counting bursts of noise even if your radio does not have a received signal indicator. If the repeater’s signal is noisy, it will be easy to tell.

2.5 Time-out timer

The repeater has a one minute time-out timer. If a signal unsquelches the receiver and the receiver remains unsquelched for one minute, the repeater sends “TO” in morse code and turns off the transmitter. When the input signal stops and the receiver squelch closes, the repeater reenables the transmitter, sends “TO” again, and normal operation resumes. One second after a signal ends, the repeater resets the timeout timer.

2.6 Touch tones

The receive audio to transmit audio connection is broken when a touch tone is detected. The connection is restored one second after the end of the tone.

3 Timing, audio level, etc.

Wait 300-500 ms after depressing the PTT (pushto-talk switch or however your transmitter is keyed) before speaking. This avoids the first syllable of your transmission being cut off. Wait about 300 ms after stopping speaking before releasing the PTT, for similar reasons. Many people have radios that turn the receivers on for 300 ms and off for 300 ms to reduce battery drain, and therefore will not hear the first 300 ms of signal transmitteed by the repeater. While it can be argued that this is their problem, pausing for slightly longer before speaking makes it more likely that the desired person will hear you.

Keep transmissions to 30 seconds or so, and definitely shorter than 1 minute. This keeps the time a person with an emergency must wait before using the repeater to a reasonable amount.

Pause for at least 1 second after another transmission ends before transmitting, in order to give someone who wishes to break in a chance (see the section on emergency traffic).

It is not necessary to wait for the repeater’s tail to drop (i.e., for the repeater to stop transmitting, which will occur 2 seconds after the previous transmission ends). The time-out timer resets after 1 second.

It is not necessary to wait for the ID to end before transmitting, as most people can understand a voice with Morse code on top of it. Often people do wait, however.

Many repeaters have a “courtesy beep” following each transmission by a second; the MIT does not. You can tell when another transmission ends by a short (50 ms?) burst of white noise (like the noise you hear when you open the squelch on your radio). It is unnecessary (and thus undesirable) to say “over” at the end of a transmission — others will hear the short burst of white noise and know that you have finished transmitting. Saying “over” is also unnecessary on simplex, as the squelch behavior on almost all FM radios is the same as the repeater’s. If you think that your signal strength is poor, it may be helpful to say “over”, but this is very rarely the case, and almost certainly more rare than most new operators think.

Speak at the proper audio level. To someone else, your voice should sound about as loud as the burst of noise as the squelch closes. This is hard to determine except from the comments of others. Ask others on the air how you sound if you aren’t sure.

You should speak clearly enough so that you may be understood even if your signal is weak. Almost everyone, myself included, could do better in this area. Speak at the proper speed. You should speak slowly enough that you are understandable, but not much more slowly. Speak concisely. If what you have to say can be gotten across in 20 seconds rather than 40, it should be.

4 Identification

FCC rules require that you identify your station at least every 10 minutes and at the end of transmissions. Good amateur practice also holds that you should identify yourself when you begin to transmit, and that you should not give your callsign excessively often. On VHF/UHF FM repeaters, this means no more frequently than once every few minutes, unless you believe that the person you are speaking to has not been able to copy your call. The repeater itself identifies every 5 minutes. It is quite reasonable to send your call every time you hear the repeater identify itself.

Do not key the PTT without speaking to see if the repeater is there. This is known as “kerchunking”, and is in violation of FCC rules requiring that transmissions be identified. Kerchunks are not properly identified, even if followed by an ID later as there is no way to tie the initial transmission to the ID. This practice is also bothersome to those responsible for maintaining the repeater because they cannot tell whether the repeater was keyed up by an amateur station making a short transmission, intermod from commercial or other transmitters, distant stations undergoing unusual propagation, a malfunction of the repeater equipment, or malicious interference.

To see if your radio and the repeater are working, transmit “N1XYZ testing”. After sending this it is completely reasonable for you not to respond to a call or to shut your radio off — you have stated that your intent was to test equipment and not to seek a QSO. Others listening will generally understand this. They may call you, but should not be upset if you do not answer.

5 Listening, calling and conversation

It is assumed throughout the examples in theis section that your call is N1XYZ.

Before transmitting, always listen to the repeater for at least 10 seconds to make sure it is not in use.

5.1 Phonetics

The word “phonetics” in the context of radio refers to using words to represent letters. This adds redundancy, enabling understanding under degraded conditions. When using phonetics, say the corresponding word rather than the letter; do not say “a as in alpha, g as in golf” — merely say “alpha golf”.

You may wish to use phonetics for your call sign at times when you have reason to believe the other person will have trouble understanding it. Such times include when you know your signal is weak or when the other person has restated your call incorrectly. It is almost never necessary to give the callsign of the other station using phonetics; an exception is when signals are weak and you wish to make sure you have copied the other station’s call correctly. It doesn’t take that much longer to use phonetics than to say letters, so use them if you find yourself having to repeat your call. Some people say “phonetics are unnecessary on FM”, but this is often not true.

The following are the ICAO (International Civil Aviation Organization) standard phonetics. You should in general use these particular words; others listening are trained to determine which of these 26 words was said given the assumption that it was one of them.

alpha bravo charlie delta echo foxtrot golf hotel india juliet kilo lima mike november oscar papa quebec romeo sierra tango uniform victor whiskey x-ray yankee zulu

If you use a random word instead, much of the benefit is lost. For instance, if I heard “ike”, I would guess that “mike” was sent, meaning “M”. If the person had sent “bike” meaning “B”, a needless error would have occurred; had they sent “bravo” and I heard “avo”, I would have correctly guessed “B”.

5.2 Listening

To indicate that you are monitoring the repeater and are willing to talk to someone, transmit “N1XYZ listening”. This is referred to as “calling listening”, as in “I called listening on the repeater at 10pm last night, and N1ABC told me my audio was distorted.” You should remain on frequency for a minute or so after sending this, in order to give someone a chance to answer you.

There is no need to announce “N1XYZ clear” if you stop listening after a minute. Thus, you should not.

Do not call listening overly frequently. Once every five minutes is definitely too frequent.

Do not make a brief contentless transmission (known as “prekerchunking”) before sending “N1XYZ” listening (to verify if the repeater is working, or other reasons), but make one transmission with your call sign and the word “listening”. If the repeater is working, you will hear it either 1) ID if it was idle or 2) send 2 seconds of silence (the repeater’s “tail”). If the repeater or your radio is not working (or, more likely, your radio is misconfigured), you will probably not hear anything. Had you sent a short contentless transmission, you would still have to identify, so there is no point in not making 1 transmission containing your callsign.

In addition, if when you “call listening” the repeater is idle, it will ID once after your call. Unless someone talks to you, it will not make any further transmissions (since it has ID’d after its last transmission). If you prekerchunk, the ID will start and the repeater will identify 5 minutes later. The repeater is often sufficently quiet that some people can actually leave a receiver on in an office environment; prekerchunking and the second ID contribute nothing useful and thus in the spirit of efficient operating practice should be avoided.

5.3 Calling

Before calling another station or calling listening, wait long enough to make sure the repeater is not in use. 10 seconds should be sufficient.

To call station N1ABC, transmit “N1ABC N1XYZ” or “N1ABC from N1XYZ”. Do not say “N1XYZ to N1ABC”. The convention is to give the other station’s call first. The reason usually stated is that the other station will hear his call even if he isn’t really listening and then listen to hear yours. Regardless of the reason, though, it is the way it is done and it confuses people to do it backwards.

Sometimes you will hear “N1ABC N1ABC N1ABC this is N1XYZ N1XYZ” or something similar. This is overly verbose, and should thus be avoided. Sending “N1ABC N1ABC from N1XYZ” is acceptable in cases when you think the called station has a radio on but may not be paying attention. You should take extra care to speak distinctly and more slowly than usual in this case, since otherwise any possible advantage of sending the other person’s callsign twice is negated by overly quick speaking.

Under normal circumstances, do not call someone overly frequently. More than twice in 15 minutes is probably too frequent. If you need to talk to someone via radio, they should know to call listening or call you when they turn their radio on.

Do not call CQ on a repeater; simply say “N1ABC listening” instead. Because of the nature of local VHF/UHF propagation and the channelized nature of FM repeater communications, someone will either not hear you even if you transmit for a minute or will hear you the first time. (On simplex, particularly a calling channel (see below) this is not true, and calling CQ is perfectly OK.)

Do not say “nothing heard N1XYZ clear” after calling someone and not receiving a reply (unless you have broken into a conversation). Each operator should be able to tell whether or not they are able to get a signal into the machine, and this transmission really has no purpose. Some busy machines have a convention that you should say this to indicate that you are done with the machine and someone else may use it, but the MIT repeater is not sufficiently busy for this to be necessary.

5.4 Answering

To answer a call from station N1ABC, transmit

1. “go ahead” or “hello fred” if you know they will recognize your voice.

2. “N1XYZ” if you know the person but think they may not recognize your voice. This has the advantage over (1) that you have identified.

3. “N1ABC N1XYZ” if you do not know the person. This is considered overly verbose by some people, but is safe in the case when you do not know the person you are working.

5.5 Fine points about calling

Because a repeater serves both to extend the range of low-power stations and as a common meeting place, it functions as a place where one can ask for help when needed. As people often listen while doing other things, being clear about whether you need help serves to keep them willing to listen and answer stations who might need help. Thus, never say “This is N1XYZ — is anyone around?” People listening will not know whether you need help or just want to talk to someone. Just saying “N1XYZ listening” indicates that you are monitoring the repeater and willing to talk.

If you do actually need help, say so. Almost anyone listening will stop what they are doing to help if you have a real emergency. For example, if the situation warrants it, you might say “This is N1XYZ with emergency traffic — is someone around to call the police to report an accident?” The word traffic here refers to radio traffic, which means roughly “a message to be sent over radio”.

If you want to demonstrate ham radio to someone else, by all means go ahead and do so. To try to find someone to talk to in order to show how the repeater works to a non-ham, say “This is N1XYZ — is anyone around for a demo?” Someone listening who is somewhat but not extremely busy may answer you, as most hams are willing to spend time to interest new people in amateur radio.

5.6 Conversation

A conversation or contact between hams is often referred to as a “QSO”.

As stated above, the repeater has a 1-minute timeout timer. After a transmission has lasted 1 minute, the repeater sends “TO” and stops transmitting. When the transmission stops, it starts transmitting again and sends “TO” again. The purpose of this is to cause users to keep their transmissions short so that someone with emergency traffic would be able to break in to use the repeater. Also, if a stray continuous signal is heard by the repeater, it will only transmit for 1 minute rather than forever.

You should in general limit transmissions to 30 seconds. If you find yourself timing the repeater out more than once in a great while, be more careful about the length of your transmissions. If you wish to talk longer than this, say “let me drop it”, stop transmitting for 1.5 seconds, and then continue. Someone who wishes to break will do so, and the person to whom you are talking will know that you intend to continue and will wait for you to do so.

If someone indicates that they have to go or otherwise indicates a desire to end the QSO, end the conversion promptly, not after talking for another minute. People often say “73”, which is a telegrapher’s code for “best wishes”. This is often used when terminating a conversation, and is always appropriate. On voice, it is pronounced “seventy-three” or “seven three”, and is not plural. Since it is a telegrapher’s abbreviation, the proper form is simply to say “seventy-three” with no attached words; do not say “I’ll say 73’s now.”

5.7 Making a call while a QSO is in progress

If two people are talking for a while and you wish to call someone on the repeater, say “N1XYZ” between their transmissions (do NOT wait the 1 second; the point of waiting 1 second is that an interrupting station does not wait and thus will be heard).

Except for emergencies, if the QSO has just begun, or will likely be over soon, wait for a bit to see if they finish before breaking in. Many QSO’s are for coordination of other activities (such as putting up antennas) and are brief.

Do not use the word “break” unless you have an emergency.

This is not a “rule” on the MIT repeater, but is on some others in the area, including the 145.23 repeater. I think it is a worthwhile convention, and suggest everyone adopt it. In addition, sending the suffix of your call conveys more information.

Do not say “call please”. This is inappropriate because the people in conversation can not tell who you are, or even if you have an amateur license. You have to ID anyway, and there is nothing gained by saying “call please” rather than your callsign.

After sending your call, wait to be acknowledged. Unless it is an emergency, your use of the repeater in the middle of someone else’s conversation should be brief.

When another station sends their call, the word “break” or words to that effect before one of your transmissions, tell them to go ahead, passing control of the repeater to them. You should do so right away rather than speaking for 30 seconds first. If they have said “break” or especially “break emergency traffic”, your next transmission should consist of ONLY “go ahead breaking station” or words to that effect.

To make a brief call to N1ABC while N1DEF and N1GHI are using the repeater: break into the conversation as above, send “N1ABC N1XYZ”. If he answers, arrange to switch to another frequency or have brief (30 seconds to a minute) words with him. Then call the person who acknowledged you when you broke in and indicate that you are through with the machine. If the person does not answer within 15 seconds or so, you should indicate “nothing heard”, ID, and pass the conversation back to the person you interrupted.

5.8 Joining an existing QSO

To join a conversation in progress, send your callsign in between transmissions, as above. Some people send only their suffix (i.e. N1XYZ might send “XYZ”); particularly if the person interrupting knows the other people and knows his suffix will be recognized. It is probably preferable to send one’s full call, but you should be aware of the suffix convention. After hearing only “N1ABC from N1DEF”, and then “N1ABC”, it is inappropriate to interrupt, as N1DEF has not yet had a chance to say what he wanted to to N1ABC. At the beginning of a QSO, wait long enough before interrupting. In general many of the same conventions that apply to interrupting in person apply on the air.

6 Character of repeaters

Different repeaters have different cultures and usage conventions. I believe that it is polite to understand and respect these conventions. This is at times problematic, since many of these conventions are unwritten, and many people wish to avoid making rules. Thus, the conventions may be not only unwritten but unstated.

6.1 Types of repeaters

There are many repeaters in the area. Monitoring a set of repeaters for a while will likely reveal to the interested listener that they are culturally different, in that what is considered appropriate use by the core users/maintainers/club members differs, at times markedly. You should attempt to understand the usage conventions of a repeater before using it heavily.

This section is incomplete; all hams are invited to submit categories and descriptions for possible inclusion.

6.1.1 Drive-time ragchew

This type of repeater is heavily used during commuting hours. There may be many people in a rotation, with everyone getting their “turn” to talk for a minute or two. Often people will turn their radios on only while commuting. It is this type of operation which gives rise to the ARRL doctrine of “base stations should yield the use of the repeater to mobile stations.”

It is considered an appropriate use of such a machine to talk on it because you are bored and have nothing better to do. However, during “prime time”, no one should monopolize the repeater, because that would deprive other commuters of their fair “turn”.

6.1.2 Semi-calling channel

This type of repeater has a regular community of users who listen frequently, but do not necessarily often transmit. Most QSOs are short (less than 5 minutes), and longer QSOs usually have a definite purpose (other than passing time).

While the regular users may call listening when they begin driving or walking, they do not view the repeater as a place to regularly hold long conversations. Users may turn to other repeaters when they wish to seek QSOs to pass time while driving.

Some of the users view the radio as an emergency tool, and often carry one or have one in the car, but do not intend to talk much. Because of the quiet character of the machine, people can actually monitor it most of the time that they are awake, even if there are non-hams around. Thus, it is a good place to look for help in an emergency.

There is usually no rule against long QSOs, but the core users rarely engage in them. There may be a sense that a long QSO once in a while is fine, but that frequent lengthy use is inappropriate. The reason for this sentiment is that frequent lengthy QSOs interfere with the use of the machine as a calling channel; while someone could break in and call, many people cannot monitor a machine with hours of conversation.

A few users, not realizing that a quiet repeater is not an ununsed repeater, might begin to use it as their “home” machine, and converse among themselves frequently and at length. This use detracts from the use of the machine as a calling channel by the regular users.

6.1.3 BARC-like

The 145.23 repeater is some “happy medium” between a ragchew machine and a semi-calling channel. Long QSOs are not discouraged, but do not happen constantly, and a significant amount of the repeater use is short QSOs. Some people do monitor often, although the machine has enough lengthy QSOs that virtually no one can listen most of the time that they are awake. Because the machine is somewhat quiet compared to some other 2m repeaters, and has many active users, this is an excellent place to look for help, and it is very likely that there will be someone near a phone to call the police.

6.1.4 Calling channel

A “calling channel” machine is one which has a (usually announced) policy that it should only be used for making contact. I do not know of any repeaters in the area with this as the announced policy.

6.2 Closed repeaters

Any repeater may be closed, in addition to being one of the above types. Being closed means that it may be used only by people who are members of the repeater. Some closed repeaters accept new members; for some membership is only by invitation.

It might seem antisocial or against the spirit of ham radio for a repeater to be closed. However, repeaters are assembled and maintained at significant expense and effort, and an open repeater may not meet the needs of the people who put it up, due to use by others that is not consistent with the intent of the machine.

There are only a few closed repeaters in the Boston area.

6.3 The MIT Repeater

The description of the “semi-calling channel” machine above fits the MIT repeater. I, for example, usually have a receiver on when awake; other people also listen a lot but do not transmit as frequently. Thus, if you “call listening” (see below) and no one answers, that does not mean that no one heard you or that no one is listening. However, if you need help, say so — people who are listening but busy will almost always respond to calls for help even if they are too busy to talk.

7 Malicous interference

At times you may hear unidentified transmissions. These may be “bootleg” operators, they may be rude comments intended to annoy but not jam, or they may be jamming. You should treat all of them the same way — by ignoring them as completely as possible. Don’t give the jammer the satisfaction he seeks by even acknowledging his presence.

7.1 Reporting jamming

Many repeaters have groups that attempt to locate jammers and persons making unidentified transmissions. To help, report the exact time, the calls of the people using the machine, the nature of the jamming, and the signal strength of the jammer. If possible, check the repeater input, and give your location, receiving equipment and antenna, and signal strength. Do not give jamming reports over the air, ever.

Send jamming reports to Greg, N1DAM, at

gdt@n1dam.lcs.mit.edu.

8 Signal reports

Giving and receiving signal reports is an important way for you and others to be able to determine how well the equipment is working and how propagation is affecting signal strengths. This section explains conventions for determining and conveying information about properties of signals.

There are three particularly important considerations determining signal quality with FM repeaters. One is the audio portion, determined by the input signal to the microphone. Another is the RF (radio frequency) path between the user’s transmitter and the antenna of the repeater; this is affected by distance between the transmitter to the repeater, among other factors. Another is the RF path between the repeater and the receiver of the other person. The RF paths are important because they determine how well the audio signal received by the microphone is reproduced at the speaker, which is of course the whole point of the radio equipment.

8.1 Audio

To understand and describe audio quality, consider conditions that cause less than ideal audio. The most common is improper level. Low audio causes the voice of the transmitting station to sound too quiet. Because whatever noise that may be present is not magically reduced when the audio level is too low, this causes the signal to noise ratio to be lowered. In addition, noise considerations aside, in order to restore the audio to an understandable level, the receiving operator must adjust his volume control (which he shouldn’t have to). Too high an audio level creates distortion caused by either clipping in either the transmitter or receiver, or by the bandwidth of the transmitted signal being too large and not fitting in either the repeater’s receiver or the receiving stations receiver. Excessive audio levels can cause squelches to close even when a signal is present.

8.2 RF

Several factors can contribute to imperfect transmission of RF signals.

low signal strength This which causes hissing noise (white gaussian noise) to appear on the demodulated audio. This can be caused by low transmitter power, a poor antenna, great distance from the repeater, intervening hills and other things. The word “interference” is not really appropriate to describe this problem, because the problem is not another signal, but lack of power in the desired signal.

mobile flutter This is really a particular form of low and varying signal strength, caused by the changing path to or from the repeater from a (usually rapidly) moving vehicle, which causes the received signal strength to fluctuate, usually periodically (with periods of a few seconds or less, depending on speed). Moving a few feet (a significant fraction of a wavelength) can often cause dramatic changes in signal strength.

interference from another station When two stations transmit on the same frequency at the same time, what a receiver hears depends on the relative signal strengths. If one is much stronger, only that one is heard. If they are comparable, audio from both is heard, along with a characteristic sound, which is hard for me to describe.

8.3 Which RF path is the problem?

In order to determine what is wrong or weak, one must determine whether it is the path from the transmitting station to the repeater or from the repeater to the receiving station that is causing the noise to be introduced. This may be done by listening to the signal from the repeater when the other station is not transmitting, during the 2-second “tail” or the ID. If the noise on the tail is equal to the noise on the other station, it is caused by the path from the repeater to you. If the tail is quiet (no noise), and the other station is noisy, it is caused by the path from the other station to the repeater.

8.4 Signal strength meters: what they don’t mean

Your radio may have an S meter. This meter measures the strength of the signal received by your radio. Thus, when you are talking to someone via a repeater, it does not indicate the strength of the other person’s transmission, and should not be given as a signal report. It does, however, indicate how well you are receiving the repeater. Be aware that S meters are not calibrated to any reference. While it is likely that the readings from your radio are consistent, there is not necessarily any fixed relationship between S9 on your radio and S9 on another radio, even if it is the same model.

8.5 Giving reports

It is important to accurately characterize a signal’s flaws in order to be able to fix them. Guidelines are given below for characterizing a signal; these are my attempt to capture what is common practice by experienced amateurs.

First, you must determine what is wrong with the signal: audio or RF, and whether it is you hearing the repeater or the repeater hearing the other person.

Audio problems are likely in the other person’s transmitter. You must of course have the volume control on your radio adjusted properly, but if their audio is distorted and the ID is not, or if their audio is much lower in level than the ID or squelch burst, the problem is almost certainly in their transmitter or voice level. Tell them, because they won’t know unless someone does.

Determine if the RF problem is “them” (path from them to repeater) or “you” (path from repeater to you). There is a third possibility, that the repeater is partially broken, but this is more difficult to determine, and until you gain more experience you should assume this is not the case. If the problem is you, do not tell the other station that he is scratchy or weak! If you are having trouble communicating, though, you can say that you are not hearing the repeater well.

If you are hearing the repeater well (tail is not noisy), then you are in a position to give a signal report to the other station. The term DFQ (dead full quieting) means that the other station’s signal is so strong that there is no hissing noise on the signal. In reality, you cannot determine what “none” is, and there will in fact never be absolutely no noise on any signal. However, none and DFQ in reality mean “so little that it sounds like none”. In order for you to be able to give this report, the repeater must be DFQ into your receiver, for otherwise you have no way of making this judgement. For the rest of the signals, the term percent quieting is used. As a signal gets stronger, the noise drops from its maximum value (when the squelch is opened on no signal, the same as the burst after someone finishes) to zero. 80% quieting means that the noise is 20% as loud as the open-channel noise. This measurement is usually not done precisely, but is guessed at by ear. It is quite likely that the entire user community of the repeater is way off in their guesses, but they are somewhat consistent.

If the signal is so weak that the repeater’s squelch is only held open part of the time, one would say, for example, “you are making it [the repeater] 70% of the time”. Signals at this level are usually understandable by experienced operators.

If a signal is fairly strong, but has periodic bursts of noise or periodic variations in signal qualtiy, this is likely “mobile flutter”. This condition should be pointed out in a signal report

You may hear people say “Q5”. This “quality” scale, or readability scale, runs from 1 to 5. 1 means not copyable, 2 barely copyable, 3 copyable with considerable difficulty, 4 copyable with some difficulty, and 5 clearly copyable (look these up and check them).

8.6 Understanding your signal strength

As you gain experience you will almost always be able to produce the correct audio level and to predict how strong your signal is into the repeater. You will remember what it is like from certain locations, power levels and antennas. Since the path conditions to the repeater are almost always analogous to the path back (not quite — different frequencies are used), you can judge your signal strength to the repeater by listening to the tail, thereby judging the strength of the repeater to you and knowing the relative abilities of the combination of your transmitter/the repeater’s receiver and your receiver/the repeater’s transmitter. The antennas and location (and thus path conditions, for the most part) will be the same, so you can use this method to guess your signal strength from new locations.

Being able to predict/determine your signal strength is important because your ability to communicate with another station is not dependent on the RF path from you to him, it is instead dependent on the path from you to the repeater and from him to the repeater — both paths must be adequately strong. This is a major feature of repeaters; if each person does their part to be able to send to/listen to the repeater, then they can communicate with each other.

9 Simplex

Simplex refers to communications without using a repeater, and is so named because the radio is set up to transmit and receive on the same frequency. Standard ARRL doctrine is that one should switch to simplex whenever possible to avoid tying up the repeater. While this may not be important, one should gain proficiency with simplex operation because 1) simplex will work if the repeater is broken or you are out of range and 2) you will gain more experience in understanding and being able to predict propagation. You may be surprised at the distances over which simplex communication works, particularly when a reasonable antenna is installed.

The national simplex frequency for the 70cm band is 446.000 MHz. This is not very active in this area, but people associated with the MIT repeater have been using it recently. While this is a calling channel, there are rarely others on it, and simplex activity is usually welcomed by them. When we start to bother others using it as a calling channel, we should pick a new frequency. We have also used 446.025 for nets and foxhunts. Either 446 or 446.025 (the latter when use will be heavy and lengthy) is suggested for the repeater community. 445.000 and 445.975 are the other two coordinated simplex frequencies for New England.

When choosing a simplex frequency, there are several considerations. First, you should be careful not to use the frequency of another repeater, particularly the input frequency. In this case you would be heard through the repeater, but would not know it! Second, for long conversations (more than a minute or so), you should not use the repeater output frequency. This is for two reasons — one, your transmissions may be “stepped on” by the repeater, and two, many people listen to the repeater often, and hearing marginal signals just causes noise in their receivers.

Often you will hear “N1ABC N1XYZ on direct”. This indicates that N1XYZ is calling N1ABC, but is doing so transmitting (simplex) on the repeater output frequency (rather than using the repeater), and would like N1XYZ to answer on simplex. This is to avoid using the repeater when it is not necessary. In general if someone calls you on simplex, you should answer the same way.

10 Touch tones

Whenever you send a touch-tone to the repeater for any reason, send your call immediately preceding it, on the same transmission. This is because tones can affect the repeater’s state, and because malicious interference often takes the form of unidentified transmissions containing tones. The control operators and trustee listen for this, and it makes their life easier to be able to tell what is going on.

11 Emergency traffic

An emergency is defined as an immediate threat to the safety of life or property, and emergency traffic is radio communications to deal with that threat.

11.1 Asking for help

When you have emergency traffic and the repeater is in use, send “break emergency traffic” immediately after the next transmission. All other stations should immediately cease transmitting (or else the station that would have transmitted next should quickly say “go ahead” to indicate to you that he heard you).

After you have control of the channel (or immediately if it was clear) give the nature of the emergency and ask precisely for the help you want. For example, “This is N1XYZ. There is a motor vehicle accident at the corner of Mass Ave. and Memorial drive in Cambridge - can someone call State Metro?”

After someone indicates they can help, they will often ask you for the location again. If they do, give it again, speaking slowly but using no unnecessary words. Stop transmitting as soon as you give the location; the responding ham may have another question. Consider the possibility that you are using a colloquial name for the location and the responding ham is not familiar with the area. Answer only the question that was asked; do not repeat the nature of the problem if the location was asked for.

For example, I once heard someone give a location as “the Southern Artery”. I questioned this, since this is not a proper name for any place. Upon questioning, the best I could do was to guess that the accident was where 93 becomes the Central Artery as the Tobin Bridge merges in (several exits north of the offramp to the airport tunnel).

11.2 Giving help

Immediately acknowledge the request for help. For example, after hearing the above call for help, one might say “This is N1XYZ. I am dialing State Metro. Please give the location again. Are there any injuries?”

The person who asked for help might then say “Cambridge. corner of Mass ave and memorial drive. can’t tell if there are injuries, but it looks bad.”

The respender would then tell the police dispatcher: “There is a motor vehicle accident at the corner of Mass Ave and Memorial drive in cambridge. Unknown if injuries, but it looks bad.”

If the accident has already been reported, they will likely say “Thank you. We already have that and we’re on the way.” Otherwise they may ask for clarification or more information. If you have to pause to get it from the other operator, say into your radio and the phone at the same time “N1ABC N1XYZ Police want to know if that is in the underpass or at street level.” N1ABC should then say “The accident is at street level.”, and the responder should repeat this into the phone.

11.3 Other aspects of emergency traffic

In general, rules may be bent or ignored during an emergency. While you may not be legally bound to follow the ham rules during a true emergency, and are not obligated to follow conventions, it is often a good idea to do so anyway, since the point of usage conventions and of communications protocols is to enable smoother communications. Thus, I recommend that you become comfortable with operating conventions before an emergency occurs; when it does they will seem natural and following them will not require thought and will make things easier.